OpenStudy (anonymous):

What are the zeroes of y = x^2 – 2x – 15? –3, –5 3, 5 –3, 5 3, –5 I believe it's D (3, -5)

OpenStudy (anonymous):

y = x^2 – 2x – 15 x^2 – 2x – 15 = 0 (x - 5) (x + 3)

OpenStudy (anonymous):

Common mistake, what is the number that will make x - 5 or x + 3 zero?

OpenStudy (anonymous):

Hint, -5 -5 = -10 it's not the right answer.

OpenStudy (anonymous):

-3, 5?

OpenStudy (anonymous):

Yeah, exactly.

OpenStudy (anonymous):

But how does that fit in with factoring?

OpenStudy (anonymous):

Look \[0 = (5-5) (5 + 3) = 0 \times 8 = 0 \] \[0 = (-3-5)(-3 + 3) = -8 \times 0 = 0\]

OpenStudy (anonymous):

I see my mistake \facepalm thanks!